# Large Language Models for Autonomous Driving (LLM4AD): Concept, Benchmark, Experiments, and Challenges

### v3

@@ V-C3 Human-in-the-Loop Validation
This subsection evaluates the effectiveness of our method in reducing takeover rates compared to the baseline system. The baseline consists of the traditional autonomous vehicle controller with default settings, where two unified controllers manage vehicle operations on longitudinal and lateral respectively. We compare this conventional approach against our VLM-based adaptive motion control system. Throughout the experiments, participants could provide explicit instructions or implicit preferences/feedback with varying degrees of directness, prompting the system to make corresponding adjustments. The instructions were categorized into three levels of directness, as defined in Tab. V.
Every participant is supposed to provide at least five instructions for each scenario. For each instruction-scenario pair, participants completed two trips - one with the baseline system and one with our VLM solution. To ensure unbiased evaluation, participants are not informed which system they are using during each trip. We use the takeover rate as our primary metric to evaluate the system’s ability to provide personalized driving experiences.
@@
V-C4 Ablation Study on Memory Module
@@
As shown in Fig. 13, With three participants, our complete system with the memory module achieves the lowest average takeover rate of 6.67%. When we remove the memory module while keeping other components the same, the average takeover rate increases substantially to 24.44%. The baseline system performs worst with a 44.44% average takeover rate. These results indicate a 72.7% reduction in takeover rate when adding the memory module to our base architecture and an 85% overall reduction compared to the baseline. This significant improvement suggests that the RAG-based memory module plays a crucial role in maintaining personalized vehicle control by effectively utilizing historical interactions and user preferences.
VI Challenges
VI-A Latency
@@
The LLM processing itself introduces substantial additional latency. Our empirical results with text-based GPT-4 [openai_gpt-4_2023] or onboard 9B VLM showed end-to-end latencies ranging from 1.2 to 1.8 seconds, while the visual-capable GPT-4o [openai2024gpt4ocard] exhibited even higher latencies around 5 seconds (assuming 300ms network latency). These latencies pose significant challenges for real-time decision-making in autonomous driving scenarios. While such response times might be acceptable for high-level decision-making or handling non-critical user requests, they become problematic for any safety-critical functions.
@@
VI-B Deployment
Another important consideration is the deployment of onboard LLMs in vehicles. While cloud-based LLM APIs offer flexibility and scalability, they rely on stable network connections, which may not always be available in all driving scenarios. To mitigate this issue, researchers are exploring the possibility of integrating LLMs directly into the vehicle’s onboard systems. However, the deployment of LLMs in autonomous vehicles faces significant computational resource limitations. Automotive hardware typically has lower computational capabilities compared to development systems, creating a substantial gap between model requirements and available resources [cui2024onboardvisionlanguagemodelspersonalized]. This constraint affects all aspects of system performance, from basic perception tasks to complex decision-making processes. The challenge is particularly obvious for LLMs, which require substantial GPU computing power that often exceeds typical automotive hardware capabilities.
Additionally, researchers can focus on optimizing LLMs specifically for deployment in vehicles. This involves streamlining the models to reduce computational requirements, enabling them to run efficiently on the limited hardware resources available in automotive systems. Model distillation, which transfers knowledge from large teacher networks, such as multimodal LLMs, to smaller student architectures, is a promising yet underexplored direction in the autonomous driving domain for addressing computational constraints [sreenivas2024llm]. Similarly, plug-and-play approaches offer another compelling avenue for integrating foundation model capabilities into on-device autonomous driving networks without adding additional computational complexity [ma2023lampilot].
Researchers must find a balance between model complexity and hardware constraints to ensure practical deployment in vehicles. As developments in hardware and model optimization continue, the feasibility of integrating powerful onboard LLMs directly into vehicles will likely increase, opening up new possibilities for enhancing the capabilities of autonomous driving systems.
VI-C Security and Privacy

### v4

@@ VI-C3 Human-in-the-Loop Validation
This subsection evaluates the effectiveness of our method in reducing takeover rates compared to the baseline system. The baseline consists of the traditional autonomous vehicle controller with default settings, where two unified controllers manage vehicle operations on longitudinal and lateral respectively. We compare this conventional approach against our VLM-based adaptive motion control system. Throughout the experiments, participants could provide explicit instructions or implicit preferences/feedback with varying degrees of directness, prompting the system to make corresponding adjustments. The instructions were categorized into three levels of directness, as defined in Tab. VII.
Every participant is supposed to provide at least five instructions for each scenario. For each instruction-scenario pair, participants completed two trips; one with the baseline system and one with our VLM solution. To ensure unbiased evaluation, participants are not informed which system they are using during each trip. We use the takeover rate as our primary metric to evaluate the system’s ability to provide personalized driving experiences.
@@
VI-C4 Ablation Study on Memory Module
@@
As shown in Fig. 15, With three participants, our complete system with the memory module achieves the lowest average takeover rate of 6.67%. When we remove the memory module while keeping other components the same, the average takeover rate increases substantially to 24.44%. The baseline system performs worst with a 44.44% average takeover rate. These results indicate a 72.7% reduction in the takeover rate when adding the memory module to our base architecture and an 85% overall reduction compared to the baseline. This significant improvement suggests that the RAG-based memory module plays a crucial role in maintaining personalized vehicle control by effectively utilizing historical interactions and user preferences.
Figure 16: This diagram shows the complete ViLaD framework for end-to-end autonomous driving: The left side illustrates multimodal input processing, where camera images and textual context are encoded and concatenated with action embeddings. The right side demonstrates the core innovation with training and inference pipelines: during training, only driving decision tokens are masked while visual/textual inputs remain unchanged; during inference, the model progressively demasks predictions starting with high-confidence decisions first. The bottom highlights ViLaD’s key advantages over autoregressive VLMs, including parallel generation, bidirectional reasoning, easy-first generation, and computational efficiency.
VII Future Trends of LLM4AD: Language Diffusion Models in Autonomous Driving
Current VLM-based end-to-end autonomous driving systems rely on autoregressive architectures that generate decisions through sequential next-token prediction. This autoregressive paradigm, while successful for many natural language processing tasks, presents some fundamental limitations when applied to the time-sensitive and spatially complex domain of real-world driving scenarios. These models generate decisions sequentially, token-by-token, which is computationally inefficient and conceptually inflexible: First, the sequential generation process creates significant inference latency, as each token must be generated one by one, making real-time decision-making challenging for safety-critical scenarios. Second, the left-to-right generation limits the model’s ability to consider global spatial relationships and future trajectory implications simultaneously, potentially leading to suboptimal driving decisions that lack comprehensive scene understanding. In the dynamic and unpredictable environment of the road, this lack of foresight and adaptability may cause a significant challenge to safety and reliability. Moreover, autoregressive VLMs suffer from the “reversal curse” phenomenon [berglund2024reversalcursellmstrained], where models trained on sequential patterns struggle with tasks requiring bidirectional reasoning or reverse inference. In autonomous driving, this limitation causes difficulty in reasoning backward from desired outcomes (e.g., “To avoid the congested traffic flow, I should exit the highway at the next exit”) or integrating future planning constraints into current decisions (e.g., “Stop at the next intersection”).
To address these inherent limitations, we propose a paradigm shift in end-to-end autonomous driving from autoregressive VLMs to the Large Vision Language Diffusion (LVLD) models [you2025lladavlargelanguagediffusion]. These models rely on a masked diffusion generation paradigm. Specifically, the text is generated by starting with a fully masked sequence, which is then iteratively filled in with tokens predicted by a model trained to reverse this masking process. Therefore, it offers several advantages in end-to-end autonomous driving:
  • Efficient Parallel Generation. Unlike sequential token prediction, diffusion models generate the decision sequences simultaneously, reducing inference latency.
  • Bidirectional Reasoning. The iterative bidirectional generation process enables models to consider both past context and future influences comprehensively.
  • Easy-First Generation Patterns. The model can adaptively focus on easier aspects of the driving task first, then progressively address more complex aspects.
With the adoption of LVLD models, we introduce ViLaD (Vision Language Diffusion), a novel end-to-end autonomous driving framework that leverages the masked diffusion generation method. This technique trains a neural network model to reconstruct a complete sequence of driving actions from a version where parts of the output are deliberately “masked.” During inference, the model begins with a fully masked (except fixed tokens) sequence of decision tokens and iteratively “unmasks” the masked tokens in parallel over several steps to produce the final output. It demonstrates a successful end-to-end implementation of the LVLD autonomous driving system from the benchmark comprehensive experiment to a real-world vehicle deployment, which is the first of its kind to the best of our knowledge.
VII-A ViLaD: Vision Language Diffusion
We present ViLaD (shown in Fig. 16), a novel architecture that leverages masked diffusion models for end-to-end autonomous driving. Our architecture consists of four key components: a vision encoder for camera-based perception, a multilayer perceptron (MLP) connector for multimodal alignment, and a diffusion-based language backbone for decision generation.
Overall Architecture Design
As shown in Fig. 16, ViLaD follows an encoder-decoder architecture where visual inputs are first processed through a vision encoder, then projected into the language embedding space through an MLP connector, concatenated with textual driving context, and finally processed by a transformer-based mask predictor to predict masks. Unlike traditional autoregressive VLMs that generate tokens sequentially, ViLaD generates complete driving action sequences through parallel diffusion-based generation, enabling real-time performance while maintaining bidirectional reasoning capabilities.
The architecture takes as input multimodal driving data: camera images $V$ and textual driving context $C$ (navigation instructions, traffic rules). The output consists of structured driving decisions $y$ representing waypoints, actions, and control commands. Mathematically, our model learns the conditional distribution $p_{\theta}(y|V,C)$ through the diffusion framework.
Vision Encoder and MLP Connector
[…]
$X_{input}=\text{Concat}([\hat{V}_{project},C_{emb}]),$ (70)
where $C_{emb}$ represents the embedded textual driving context.
Diffusion Language Backbone
The core of ViLaD is a masked diffusion transformer that processes the concatenated multimodal tokens through bidirectional self-attention, enabling joint reasoning over visual and textual information.
Fix Pattern Training. The backbone employs a bidirectional transformer without causal masking, processing the concatenated multimodal sequence through transformer layers. The bidirectional self-attention enables the model to consider the complete context from both visual and textual modalities when predicting individual driving decisions.
During training, we utilize the masked diffusion approach specifically adapted for multimodal conditional generation and end-to-end autonomous driving tasks. Specifically, our input is a multimodal input $X_{input}$ consisting of projected visual features $\hat{V}_{project}$ and driving context embedding $C_{emb}$. The target output $y_{0}$ is a fixed-pattern driving sequence that includes both control actions and punctuation marks, where punctuation marks always appear in predefined positions and the length of the $y_{0}$ is always consistent. This design ensures the model focuses solely on predicting action values rather than unnecessary formatting. We construct the model input as $X_{t}=\text{Concat}([X_{input},y_{0}])$, applying masks only to the action tokens in $y_{0}$ while keeping $X_{input}$ and the punctuation tokens visible throughout training.
TABLE XII: End-to-end planning performance on the nuScenes dataset (* results from  [xing2025openemma]).
Method Learning L2 (m) 1 s ($\downarrow$) L2 (m) 2 s ($\downarrow$) L2 (m) 3 s ($\downarrow$) L2 (m) Avg ($\downarrow$) Failure Rate (%) ($\downarrow$)
LLaVA-1.6-7B* Zero-Shot 1.66 3.54 4.54 3.24 4.06
Llama-3.2-11B* 1.50 3.44 4.04 3.00 23.92
Qwen2-VL-7B* 1.22 2.94 3.21 2.46 24.00
ViLaD-Zero 1.09 2.51 3.74 2.45 0.03
LLaVA-1.6-7B* Open-EMMA 1.49 3.38 4.09 2.98 6.12
Llama-3.2-11B* 1.54 3.31 3.91 2.92 22.00
Qwen2-VL-7B* 1.45 3.21 3.76 2.81 16.11
ViLaD-CoT 1.24 2.74 4.13 2.71 0.17
LLaVA-1.6-7B SFT 0.91 2.50 3.44 2.28 55.25
Llama-3.2-11B 0.80 2.31 3.10 2.07 0.06
Qwen2-VL-7B 1.32 2.94 3.98 2.74 0.03
ViLaD-SFT 0.79 1.92 2.83 1.85 0.00
ViLaD-Opt Efficient Optimized SFT 0.81 1.93 2.69 1.81 0.00
In each training step, for a training sample $X_{t}=\text{Concat}([X_{input},y_{0}])$, we sample a noise level $t\sim U(0,1)$ and independently mask each token in the target action tokens with probability t:
$q(y_{t}|y_{0})=\prod_{i=1}^{L}q(y_{t}^{i}|y_{0}^{i}),$ (71)
where $L$ is the length of the driving sequence $y_{0}$. The projected visual features $\hat{V}_{project}$, textual context $C$, and punctuation marks remain unmasked throughout the process.
The training objective is:
$\mathcal{L}(\theta)=-\frac{1}{t}\sum_{i=1}^{L}\mathbb{1}[y_{t}^{i}=\text{[M]}]\log p_{\theta}(y_{0}^{i}|X_{t}),$ (72)
where $\mathbb{1}[y_{t}^{i}=\text{[M]}]$ ensures the loss is computed only on masked driving decision tokens, and $X_{t}$ represents the full concatenated sequence with masked driving responses.
Algorithm 1 Remasking Process
1: Current masked sequence $y_{t}$, predictions $\hat{y}_{0}$, total number of tokens $L$, total steps SpsSps, threshold $\tau$
2: Next sequence state $y_{s}$
3: $\textit{confidence}_{i}\leftarrow p_{\theta}(\hat{y}_{0}^{i}\mid X_{t})$ for $i=1,\dots,L$
4: $\text{masked\_indices}\leftarrow\{i\mid y_{t_{i}}=\text{[M]}\}$
5: $n_{\text{unmask\_schedule}}\leftarrow\lceil\frac{L}{Sps}\rceil$ $\triangleright$ Number of tokens to unmask based on schedule
6: $n_{\text{unmask\_threshold}}\leftarrow|\{i\in\text{masked\_indices}\mid\textit{confidence}_{i}>\tau\}|$ $\triangleright$ Number of tokens to unmask based on confidence
7: $n_{\text{unmask}}\leftarrow\max(n_{\text{unmask\_schedule}},n_{\text{unmask\_threshold}})$
8: $\text{conf}_{\text{masked}}\leftarrow\{i\in\text{masked\_indices}|\textit{confidence}_{i}\in\text{topk}(\textit{confidence},n_{\text{unmask}})\}$ $\triangleright$ Get top-n confident tokens
9: for $i\in\text{conf}_{\text{masked}}$ do
10:   $y_{s}^{i}\leftarrow\hat{y}_{0}^{i}$ $\triangleright$ Unmask the selected predictions
Fix Pattern Confidence-based Inference. During inference, we perform the reverse diffusion process starting from a partially masked sequence $y_{1}$, where all target action tokens are masked while punctuation tokens and length remain fixed, consistent with the training setup. The model then iteratively refines the masked tokens over up to $Sps$ denoising steps to generate the final driving action sequence.
One primary contribution we achieve is focusing on addressing quality degradation in parallel token generation caused by the conditional independence assumption in diffusion LLMs. When unmasking multiple token positions $i$ and $j$, MDMs sample these from $p_{\theta}(y^{i}_{0}|y_{t})\cdot p_{\theta}(y^{j}_{0}|y_{t})$ due to the conditional independence assumption, while the true joint probability is $p_{\theta}(y^{i}_{0},y^{j}_{0}|y_{t})=p_{\theta}(y^{i}_{0}|y_{t})\cdot p_{\theta}(y^{j}_{0}|y_{t},y^{i}_{0})$. This discrepancy degrades generation quality. To address this, enlightened by Fast-dLLM [wu2025fastdllmtrainingfreeaccelerationdiffusion], we propose confidence-aware parallel decoding: at each iteration, we compute confidence scores for each token and only unmask those exceeding the threshold $\tau$. If the number of tokens with confidence scores exceeding $\tau$ is less than $\lceil L/Sps\rceil$, we unmask the $\lceil L/Sps\rceil$ highest-confidence tokens to ensure progress.
The inference steps begin with sampling. At a step with the current sequence $X_{t}$, we use the model to predict all masked tokens simultaneously:
$\hat{y}_{0}^{i}=\underset{y_{0}^{i}}{\arg\max}\,p_{\theta}(y_{0}^{i}\mid X_{t})\quad\text{for all }i.$ (73)
Then, following the threshold-based remasking strategy, we keep at least $\lceil L/Sps\rceil$ high-confidence predictions unchanged, and remask all the other tokens. The remask process is summarized in Algorithm 1. After remasking, the resulting sequence is used as the input sequence in the next step. This iterative process lasts for at most $Sps$ steps until all tokens are masked, following the standard discretized reverse diffusion schedule. Both the number of sampling steps $Sps$ and the threshold $\tau$ can be adjusted to balance between generation quality and inference speed. The whole inference process is summarized in Algorithm 2.
Algorithm 2 ViLaD Inference
1: Visual features $\hat{V}$, driving context embedding $C_{emb}$, driving sequence length $L$, steps SpsSps
2: Predicted driving sequence $y_{0}$
3: Initialize: $X_{1}\leftarrow\text{Concat}([\text{MLP}(\hat{V}),C_{emb},[\text{[M]}]^{L}])$
4: while $y_{s}$ contains [M] do $\triangleright$ Loop until all tokens are unmasked
5:   $\hat{y}_{0}\leftarrow\arg\max_{y_{0}}p_{\theta}(y_{0}|X_{t})$ $\triangleright$ Parallel prediction
6:   $y_{s}\leftarrow\text{Remask}(y_{t},\hat{y}_{0},p_{\theta}(\hat{y}_{0}^{i}|X_{t}))$ $\triangleright$ Remask Process
7:   $X_{t}\leftarrow\text{Concat}([V_{project},C_{emb},y_{s}])$
8: return $y_{0}$
VII-B Experiment and Results
VII-B1 Experiment Setup
We conduct experiments on the nuScenes dataset [caesar_nuscenes_2020] using only front camera images from the last frame and ground-truth trajectories for training and validation. Evaluation focuses on L2 error in the first 3 seconds and inference latency for real-time performance. Following Open-EMMA [xing2025openemma], a prediction is considered a failure if the L2 distance exceeds 10 within the first second of the future trajectory. We compare ViLaD against three state-of-the-art VLMs: LLaVA-1.6-Mistral-7B [li_llava-med_2023], Llama-3.2-11B-Vision-Instruct [zhang_video-llama_2023], and Qwen2-VL-7B-Instruct [wang2024qwen2vlenhancingvisionlanguagemodels] under zero-shot, OpenEMMA prompting [xing2025openemma], and supervised fine-tuning (SFT) settings.
We evaluate multiple configurations of our proposed ViLaD framework to analyze different aspects of performance and capabilities. We utilize a pretrained LLaDA-V [you2025lladavlargelanguagediffusion] model as our base model, and fine-tune from it. Other ViLaD variants include ViLaD-SFT (full model with diffusion-based SFT), ViLaD-Zero (zero-shot), ViLaD-CoT (chain-of-thought reasoning using OpenEMMA [xing2025openemma]), and ViLaD-Opt (optimized inference), offering a broad evaluation of performance and capabilities.
VII-B2 Optimization Strategy
The outcome from dLLM-Cache [liu2025dllmcacheacceleratingdiffusionlarge] is used in our work. dLLM-Cache [liu2025dllmcacheacceleratingdiffusionlarge] introduces a training-free adaptive caching framework that combines long-interval prompt caching with partial response updates guided by feature similarity. These methods accelerate the inference speed of ViLaD and hence reduce the latency of the whole framework.
Figure 17: The comparison of inference time between VLMs and ViLaD-Opt.
TABLE XIII: On-Board Model Performance Metrics for Interactive Parking.
Scenario Model Performance Safety Metrics Comfort Metrics
Latency (s)
Success Rate (%)
Long. Vel. Var.( m²/s²)
Lat. Vel. Var.($10^{-4}$ m²/s²)
Max Long. Accel.(m/s²)
Max Lat. Accel.(m/s²)
Parking Selection 0.18 95.18 1.07 2.66 1.43 1.12
E2E Instruction Following 0.32 85.91 0.62 1.53 1.39 0.68
VII-B3 End-to-End Planning
Tab. XII presents comprehensive comparison results across different methods. Our ViLaD framework demonstrates the best performance across all evaluation metrics compared to baseline VLM approaches. In the zero-shot setting, ViLaD-Zero achieves remarkable performance with L2 errors of 1.09 m, 2.51 m, and 3.74 m at 1 s, 2 s, and 3 s horizons, respectively, significantly outperforming the best baseline (Qwen2-VL-7B [wang2024qwen2vlenhancingvisionlanguagemodels]), which achieves 1.22 m, 2.94 m, and 3.21 m. Most notably, ViLaD-Zero maintains an extremely low failure rate of 0.03% compared to baseline failure rates ranging from 4.06% to 24.00%, demonstrating the inherent robustness of our diffusion-based approach.
The OpenEMMA prompting framework shows consistent advantages of our approach. ViLaD-CoT achieves competitive performance with an average L2 error of 2.71 m while maintaining a near-zero failure rate (0.17%), outperforming all baseline methods in this category. However, we found that the ViLaD model shows no improvement using the chain-of-thought prompting method. The chain-of-thought prompting method still needed to be explored.
With only the SFT setting (no fixed pattern training and fixed pattern confidence threshold inference), ViLaD-SFT achieves the best overall performance with L2 errors of 0.79 m, 1.92 m, and 2.83 m across the three time horizons, resulting in an average L2 error of 1.85 m with zero failure rate. This represents significant improvements over the best baseline (Llama-3.2-11B [zhang_video-llama_2023] SFT), which achieves 2.07 m average L2 error. Notably, while some baseline methods achieve competitive L2 errors, they suffer from substantially higher failure rates (up to 55.25% for LLaVA-1.6-7B [li_llava-med_2023] SFT), highlighting the safety advantages of our approach.
ViLaD-Opt, our optimized deployment configuration with fixed pattern training and fixed pattern confidence threshold (0.5) inference, maintains the most competitive performance (1.81 m average L2 error) while achieving inference speedup, making it suitable for autonomous driving applications.
VII-B4 Inference Efficiency Experiment
To validate the real-time performance of our framework, we conducted a comprehensive inference efficiency experiment, comparing our approach against leading autoregressive VLM baselines that benefit from highly mature optimization technologies. The VLM baselines (LLaVA-1.6-7B [li_llava-med_2023], Qwen2-VL-7B [wang2024qwen2vlenhancingvisionlanguagemodels], and Llama-3.2-11B [zhang2023videollamainstructiontunedaudiovisuallanguage]) are accelerated using standard KV caching, while our ViLaD-Opt model, representing a still-developing class of vision language diffusion models, utilizes a dLLM cache. The results from the experiment, conducted on a single NVIDIA A100 GPU and presented in Fig. 17, show the significant computational advantages of our diffusion-based method. Without caching optimizations, ViLaD-Opt achieves an inference time of 4.28 s, proving to be 6.5 to 14 times faster than the autoregressive models, which require between 28.10 s and 60.18 s. While caching accelerates all models, ViLaD-Opt maintains its superiority with an inference time of 2.78 s, still outperforming the fastest baseline (3.18 s). This experiment empirically confirms that the parallel generation paradigm of ViLaD provides a critical advantage in inference speed by avoiding the token-by-token processing bottleneck in autoregressive architectures, making it highly suitable for deployment in safety-critical, real-time applications.
VII-C Real-World Case Studies: On-Board Interactive Parking
We further validate the real-world applicability of our approach through two on-board case studies (interactive parking and instruction-following driving). Both of them deployed on the same vehicle platform as shown in Fig. 8.
VII-C1 System Setup
We evaluate system efficiency (inference latency) and driving quality (success rate, safety, and comfort). A trial is successful if the executed driving decision matches the instruction. Safety is measured by the variance of longitudinal and lateral velocities, and comfort by maximum longitudinal/lateral accelerations.
Two lightweight models fine-tuned from SMDM [nie2025scalingmaskeddiffusionmodels], which are lightweight and support more on-board inference compared with the larger-scale models used in the full framework above, are used for the respective tasks. The parking dataset contains $\sim$3,000 command–action pairs covering diverse slot preferences, and the instruction-following dataset has a similar scale focused on short-horizon commands. Each task is tested with 300 unseen instructions. Demo videos are publicly available. [^5] Real-vehicle experiment demo video with ViLaD: https://youtu.be/pNfC7wrENyk
VII-C2 Case Study I Interactive Parking
[…]
VII-C3 Case Study II End-to-End Instruction Following
For open-road commands (e.g., “Turn left and stop near the crosswalk”), the vehicle autonomously executes the instruction over 28–38 m. ViLaD reaches an 85.91% success rate with 0.32 s latency, demonstrating real-time performance. Stable motion (velocity variances: $0.62\text{\ }{\mathrm{m}}^{2}\text{\,}{\mathrm{s}}^{-2}$ / $1.53\text{\times}{10}^{-4}\text{\ }{\mathrm{m}}^{2}\text{\,}{\mathrm{s}}^{-2}$) and low accelerations ($1.39\text{\ }\mathrm{m}\text{\,}{\mathrm{s}}^{-2}$, $0.68\text{\ }\mathrm{m}\text{\,}{\mathrm{s}}^{-2}$) show reliable control and comfort. These results highlight ViLaD’s ability to generalize from parking to dynamic, instruction-driven driving.
VIII Challenges
VIII-A Latency
@@
The LLM processing itself introduces substantial additional latency. Our empirical results with text-based GPT-4 [openai_gpt-4_2023] or on-board 8B VLM showed end-to-end latencies ranging from 1.2 to 1.8 seconds, while the visual-capable GPT-4o [openai2024gpt4ocard] exhibited even higher latencies around 5 seconds (assuming 300ms network latency). These latencies pose significant challenges for real-time decision-making in autonomous driving scenarios. While such response times might be acceptable for high-level decision-making or handling non-critical user requests, they become problematic for any safety-critical functions.
@@
VIII-B Deployment
Another important consideration is the deployment of on-board LLMs in vehicles. While cloud-based LLM APIs offer flexibility and scalability, they rely on stable network connections, which may not always be available in all driving scenarios. To mitigate this issue, researchers are exploring the possibility of integrating LLMs directly into the vehicle’s on-board systems. However, the deployment of LLMs in autonomous vehicles faces significant computational resource limitations. Automotive hardware typically has lower computational capabilities compared to development systems, creating a substantial gap between model requirements and available resources [cui2024onboardvisionlanguagemodelspersonalized]. This constraint affects all aspects of system performance, from basic perception tasks to complex decision-making processes. The challenge is particularly obvious for LLMs, which require substantial GPU computing power that often exceeds typical automotive hardware capabilities.
Additionally, researchers can focus on optimizing LLMs specifically for in-vehicle deployment. This involves streamlining the models to reduce computational requirements, enabling them to run efficiently on the limited hardware resources available in automotive systems. Model distillation, which transfers knowledge from large teacher networks, such as multimodal LLMs, to smaller student architectures, is a promising yet underexplored direction in the autonomous driving domain for addressing computational constraints [sreenivas2024llm]. Similarly, plug-and-play approaches offer another compelling avenue for integrating foundation model capabilities into on-device autonomous driving networks without adding additional computational complexity [ma2023lampilot].
Researchers must find a balance between model complexity and hardware constraints to ensure practical deployment in vehicles. As developments in hardware and model optimization continue, the feasibility of integrating powerful on-board LLMs directly into vehicles will likely increase, opening up new possibilities for enhancing the capabilities of autonomous driving systems.
VIII-C Security and Privacy